LUCI/LBC Technical Observing, prep for Partner Science:
UT 2011 Feb 25

Observer: M.Pedani
Telescope Operator: S.Allanson
Instrument Support: T.Shih
Telescope Support: J.Urban
SW support: D.Cox

Summary

After the mask Exchange in the afternoon, made some Functional Checks of LUCI1. After a bumpy start we are up and running smoothly.
Everything seems to run fine.
We checked/tested the following:

  • Mask Tables after exchange
  • CAL Unit IN/OUT & LAMPS ON/OFF
  • Focus Gradients both with 3.75 and 1.8 cameras (Sieve mask)
  • LUCI1 Flexure Test
  • Run science-type dice5 pattern
  • Test defocus on AGw by 13mm, Johanna's acq+sci script
  • Biases with LBCB
  • Test the new DOFPIA version
  • Verified that a Red_Master OBs can be played with LBCB only and that LBCB can guide the mount and can send z4 correction during Sci Exposures
  • LBC Collimation; we had to give by hand +12000nm Z4 and +1200nm Z7 corrections at the begin of the session!
  • Taken some Field Check data for LUCI until we closed due to high winds

Details

Mask Exchange done by Tim in the afternoon.

23:40 LUCI put back from SAFE to OPERATIONAL mode
00:15 LUCI at 77K

LUCI1 functional checkout:

showSupervisorStatus shows all OK
Web-IO shows power on to MCE/ROE and CAL Unit ON
UI Startup OK, Initialize instrument OK, Moved Blind Mask to FPU OK
Error while connecting to GEIRS server, closing everything and re-starting GEIRS server
Second attempt OK
One image saved to check archiving; OK
Check system time vs our GPS; OK
00:53 Moving Calibration Unit IN. Taking more than 10 min; something wrong here but no error message
01:06 Issuing again CAL UNIT IN and getting same message "Moving Calibration Unit IN"
01:15 Stop CAL UNIT service and restart; also reinit LUCI; now OK
01:19 Moving CAL Unt IN; OK
01:20 Switch ALL LAMPS ON and OFF; OK
01:22 Moving CAL Unit OUT; nothing seems to happen again. I saw the yellos message "moving" after trying 5-6 times. Eventually it moved OUT
01:26 Try to set-up a new DIT and get Error message "2011.02.25 01:26:10 s Connection to GEIRS Server timed out while waiting for answer!
01:30 Turn OFF LUCI Rack Electronics and ON again
01:36 Restart GEIRS server
01:37 Restart LUCI UI and get all Kind of errors one of them is critical: "Critical level reached for sensor rack temp 2: 289.63 K, 1.43 K/min"

01:40 Detector Temp is now 70.958K instead of 77K ! Tim put it back to OPERATION mode and T start rising again; same problems unable to initialize almost every Unit
01:55 Restart ALL LUCI software
02:30 We are up and running

02:35 Check mask tables after exchange, OK

02:45 Adjust Pointing IA=-90 CA=70

02:50 Check darks O2DCR 2.0s (11.75 e-) OK. Images 0013,0014
03:00 Check darks MER 10s (5.51e-) OK. Images: 0015,0016

03:05 Check Focus gradients. Optical Sieve Mask & 3.75 camera. Image: 0017, J 20s. Looks OK
03:15 Check Focus gradients. Optical Sieve Mask & 1.80 camera. Image: 0018, J 20s. Look OK but the probe was on axis....
03:23 Check Pointing Again. IE=-105 CA=78; Note that IE changed by -15" since 2:23 UT, CA changed by +8"
03:29 Check Focus gradients. Optical Sieve Mask & 3.75 camera. Image: 0019, J 40s. (now probe is out)
03:34 Check Focus gradients. Optical Sieve Mask & 1.80 camera. Image: 0020, J 40s. (now probe is out). FWHM ~ 2 pix at top, ~1.5 at bottom center

03:43 Weather Conditions: T=-6.2 T_mirror = -4.8 RH= 35 Wind=11 m/s. Sky Clear
03:44 Slew to BS9142 to check Collimation. We go to SE to avoid some wind. Seeing is ~1.5" - 2.5" highly variable now

04:16 Run LUCI1 FLEXURE TEST (script is /LBTO/Flexures/RotatorScript.txt) Images: 0023 - 0052

DJT 01MAR2011 - Data not valid for this test...
   Observations taken on-sky in POSITION/ACTIVE mode!

04:34 Recollimate on BS9142. Seeing is now ~1.5 - 2.6" ! We are gonna wait in this position for a while. We saw up to 4.5" seeing

05:15 Weather Conditions: T=-5.5 T_mirror=-4.9 RH=35% Wind = 9 m/s. Seeing is just crazy from 0.8" to 4.0" !!
05:25 We closed ALL the Vent Doors. We have pretty decent seeing between 1.1" and 2.0" so we go for the next test

05:28 Run science-type dice5 pattern on BS9142. Images:0053-0102

DJT 01MAR2011 - script set up with REL DETXY offsets but it should be ABS.  
Script changed.  Data analyzed as taken...evidence for 0.23 degree relative 
rotation between detector coordinates and AGw transform.  RMS of the pattern 
is ~0.5 pixels (60mas), a little high.
### Geomap output:
Coordinate mapping status
    Xin and Yin fit rms: 0.6407147  0.5049252
Coordinate mapping parameters
    Mean Xref and Yref: 301.7786  498.8227
    Mean Xin and Yin: 300.  500.
    X and Y shift: -1.332839  -1.309814  (xin  yin)
    X and Y scale: 1.004495  1.002512  (xin / xref  yin / yref)
    X and Y axis rotation: 359.77518  359.77518  (degrees  degrees)

06:36 Weather Conditions: T=-5.6 T_mirror=-4.7 RH=43 Wind=10m/s Seeing ~ 1.1"

06:35 Test defocus on AGw by 13mm. Change Config File. Focus is now 44 instead of 31 (default)

06:46 Slewing to Field WASP12 to do the Test above; WFS pupils array is missing the lower right part (vignetting)

07:07 Starting science script on WASP12b (2" slit); seeing is now ~2.5". Images: 0113-0128

DJT 01MAR2011 - Test appears to be successful.  

07:48 LUCI test Done for the moment. LUCIFER look OK for Sci Ops

07:48 Weather Conditions: T=-4.8 T_mirror=-4.8 RH=35% Wind=12 m/s. Sky is still clear

07:48 We Switch to LBCB; in the meantime LBCB had been warmed-up and turned on successfully

07:56 Taking 12 Biases with LBCB while Steve is switching to LBCB. Images: lbcb.20110225.075613 to 080226. Median Bias level=827 ADU on Chip #2. They look OK

08:15 LBC crashes every time we connect to LBT. We realize that we did not switch to Build AO5. So we do that

08:23 Slew to M101 Focus Field

08:26 Running DOFPIA to test the NEW DOFPIA Code

RB_rVfastextra is not loaded by the LBC UI and we get this error:

Waiting for OB to finish and data files to appear in /newdata % Compiled module: WAIT4OBFILE_BLUE. % Compiled module: GET_DATE. % Compiled module: DAYCNV. ***** WARNING ***** Data files DID NOT appear in /newdata Pls double check and then: press Q to exit or any other key to continue

08:31 Loading RB_rVfastextra manually to take one shot of LBCB Pupil; Image: 083120

08:32 Try again DOPFIA, /blueonly. Again Nothing happened. The OB is not loaded by DOFPIA. The Red Channel and Mirror boxes are unticked

08:41 try to restore the previous version of DOFPIA. In /home/lbcobs/LBCFPIA/lbcfpia the /src directory is actually the latest version. I rename it to src-LATEST and rename the previous one (src-20110218/) back to src

08:45 Try the Old Code. DOPFIA, /blueonly. It's working, the B_Vfastextra OB is loaded and executed. So it looks like the new code has a bug

Computed Aberrations (nm) | (arcsec) Defoc AstX AstY ComaX ComaY Sph3 Sph5 | Est. Est. (Z4) (Z5) (Z6) (Z7) (Z8) (Z11) (Z22)| seeing WIQ

-15852 0 0 0 0 0 0 2.32 6.25 lbcb.20110225.084746.fits

Then no pupils found...aborted OB

08:53 Running again the old DOFPIA, no good pupil found but it worked.
09:08 Retry with the latest DOFPIA. Same result the OB is not played. On the phone with Olga

09:14 Playing the B_Vfastextra OB manually and it works: Image 091428

09:15 Olga changes some permissions. (Note added later by OPK: Files as downloaded from svn had only read permission for group & world (rwxr--r--). I ran "chmod 755 *". This probably did not matter for the *pro IDL routines, but it did for the *pl perl files which should load and play the focus OB; someone from the software group could confirm.) We give manually z7=+1200nm and z4=12000nm

09:20 Try again the latest DOFPIA version and it seems to work now!

lbcb.20110225.092033.fits -616 -1097 -714 2039 -988 0 0 2.04 1.12

lbcb.20110225.092154.fits 1006 -128 770 -620 426 -429 -39 1.35 0.64

lbcb.20110225.092311.fits -627 99 -447 121 253 -64 89 1.09 0.32

lbcb.20110225.092429.fits 753 -156 355 95 164 -285 28 1.20 0.38

lbcb.20110225.092545.fits 57 452 -475 -247 -303 -249 79 1.00 0.32

lbcb.20110225.092704.fits -181 -427 -25 -235 -87 2 77 1.56 0.20

lbcb.20110225.092823.fits 717 -365 -77 182 170 -335 11 1.20 0.39

lbcb.20110225.092939.fits 671 808 24 -10 -23 -343 -4 1.09 0.42

lbcb.20110225.093054.fits -467 -40 249 -101 41 -46 7 1.28 0.22

lbcb.20110225.093210.fits 261 -286 41 -58 150 -163 -9 1.33 0.20

It converged after 10 iterations !!

09:38 Slew to WT10_376 to Check Pointing close to M82; IE=-37.36, CA=-145.02

09:42 Slew to M82 to Collimate

09:44 Run dofpia,/blueonly

lbcb.20110225.094516.fits 61 273 -391 87 79 128 100 1.38 0.19

lbcb.20110225.094633.fits -130 127 -331 -90 125 75 81 1.25 0.16

lbcb.20110225.094751.fits -49 44 184 -188 -4 -107 -7 1.11 0.15

Blue Convergence reached for seeing of 1.11279 arcsec. So 3 iteration to converge this time

09:50 Now we test on M82 how the Red_master OB is run if we only have LBCB. OB is : M82_BRSY_Rmaster.ob

We can CONFIRM that even with a Red-Master OB, if we untick the Red camera and Run the OB, LBC recognizes that there is only the Blue Camera and it considered as MASTER. In fact, the Blue Camera is sending Offset to the Mount:

grep arcs /lbt/log/2011/02/20110225.log

Feb 25 09:53:35 tcs2 LBT_PCS: Centroid Xrad: -0.000000196554042 Yrad: 0.000000896431145 Xarcs: -0.040542181478162 Yarcs: 0.184902196377498
Feb 25 09:53:53 tcs2 LBT_PCS: Centroid Xrad: 0.000000029595812 Yrad: -0.000000556750972 Xarcs: 0.006104574413138 Yarcs: -0.114838131293233
Feb 25 09:54:16 tcs2 LBT_PCS: Centroid Xrad: -0.000000189219600 Yrad: 0.000000434617155 Xarcs: -0.039029344133156 Yarcs: 0.089646223332037
Feb 25 09:54:33 tcs2 LBT_PCS: Centroid Xrad: -0.000000588646436 Yrad: 0.000000117997271 Xarcs: -0.121417043083049 Yarcs: 0.024338684257456
Feb 25 09:54:42 tcs2 LBT_PCS: Centroid Xrad: 0.000000502106067 Yrad: -0.000000267969944 Xarcs: 0.103566810617982 Yarcs: -0.055272768628676
Feb 25 09:54:50 tcs2 LBT_PCS: Centroid Xrad: -0.000000259534327 Yrad: 0.000000371710821 Xarcs: -0.053532797754271 Yarcs: 0.076670860537476
Feb 25 09:56:11 tcs2 LBT_PCS: Centroid Xrad: 0.000000000000000 Yrad: 0.000000000000000 Xarcs: 0.000000000000000 Yarcs: 0.000000000000000
Feb 25 09:56:20 tcs2 LBT_PCS: Centroid Xrad: -0.000000064485737 Yrad: 0.000000477685597 Xarcs: -0.013301138090510 Yarcs: 0.098529727200728
Feb 25 09:56:36 tcs2 LBT_PCS: Centroid Xrad: -0.000000477378343 Yrad: 0.000000004813169 Xarcs: -0.098466351446613 Yarcs: 0.000992787280778
Feb 25 09:56:53 tcs2 LBT_PCS: Centroid Xrad: -0.000000026687911 Yrad: -0.000000551889947 Xarcs: -0.005504776704804 Yarcs: -0.113835472908862<br>

Also NO Z2,Z3 (Tip/Tilt) corrections are being sent to the Blue mirror and z4 corrections are being sent to the Blue mirror to keep the star in focus since we have 1 star in Tech Chip #2

10:16 M82 OB finished successfully; Images are: lbcb.20110225.095314 to lbcb.20110225.101543.fits

10:35 Weather Conditions. Wind speed is picking up now T=-4.0 T_mirror=-4.0 RH=35% Wind up to 19.5m/s!

10:39 Switch back to BP7 and LUCIFER but we close dome for safety given the wind speed

11:29 Back in business. Slew to BS9162 to adjust Pointing; IE=-96.0, CA=70.0

11:44 Slew to BS9164 for some Field Check Data Collection before twilight

11:49 field_check, /GCS and pick up BS9164_short_field.list

First Image: left_wfscimage000270.fits

Script crashes soon:

=> Running collimate_process: /Repository/GCSfiles/dailyimages//left_wfscimage000272.fits ===> Error in gcs_read_info: file not found

Last image taken: 0275

12:01 Restart script: field_check, /GCS, START=1

First Image: 0285

Script crashes again:

===> Error in irc_presetTelescope err = Resmsg[0] = PresetTelescope result status: Error err = Resmsg[1] = no guide star found in guide box after resumeGuiding % Program caused arithmetic error: Floating illegal operand

12:07 Restart GCS and Restart script: field_check, /GCS, START=1

Images: 0286 to 311 (script stopped due to high wind)

12:38 Wind reaches limits and we close. We call it a night

-- MarcoPedani - 24 Feb 2011
Topic revision: r4 - 01 Mar 2011, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback